A 500 g bar of ice is initially at -25 °C. Calculate the amount of energy required to vapourize 25% of the substance at 100 °C


1
Expert's answer
2021-06-01T10:47:50-0400

Specific heat capacity for Ice = 2.1 J/g°C

Latent heat of fusion of Ice = 335 J/g

Specific heat capacity for Water = 4.2 J/g°C

Latent heat of evaporation: 2256 J/g

"Q_1 = 25\times 500 \times 2.1 = 26250 \;J" (from -25 °C to 0 °C)

"Q_2 = 100 \times 500 \times 4.2 = 210000 \;J" (from 0 °C to 100 °C)

"Q_3 = 500 \times 335 = 167500 \;J" (heat of melting)

"Q_4 = 500 \times 2256 = 1128000 \;J" (heat of evaporation)

Total heat "Q_T = 26250+210000+167500+1128000= 1531750 \;J"

"Q_{0.25}=1531750 \times 0.25=382937.5 \;J = 382.93 \;kJ"

Answer: 382.93 kJ